OpenStudy (awy):

When a circular plate of metal is heated in an oven, its radius increases at a rate of 0.04 cm/min. At what rate is the plate's area increasing when the radius is 40cm?

OpenStudy (anonymous):

\[A = \pi*r^2\] \[dA/dt = 2*\pi*r*(dr/dt)\]

OpenStudy (anonymous):

you got it ...

OpenStudy (awy):

so dr/dt = .00157 cm/min. This helped, ty.